### v2

# Geometric Framework for Cell Oversegmentation

Abstract
3D cell segmentation methods are often hindered by oversegmentation, where a single cell is incorrectly split into multiple fragments. This degrades the final segmentation quality and is notoriously difficult to resolve, as oversegmentation errors often resemble natural gaps between adjacent cells. Our work makes two key contributions. First, for 3D cell segmentation, we are the first work to formulate oversegmentation as a concrete problem and propose a geometric framework to identify and correct these errors. Our approach builds a pre-trained classifier using both 2D geometric and 3D topological features extracted from flawed 3D segmentation results. Second, we introduce a novel metric, Geo-Wasserstein divergence, to quantify changes in 2D geometries. This captures the evolving trends in cell mask shape changes in a geometry-aware manner. We validate our method through extensive experiments on in-domain plant datasets, including both synthesized and real cases, as well as on out-of-domain animal datasets to demonstrate transfer learning performance. An ablation study further highlights the contribution of the Geo-Wasserstein divergence. A clear pipeline is provided for end-users to build pre-trained models to any labeled dataset.
1 Introduction
3D cell segmentation aims to reconstruct complete 3D cellular structures from sequential 2D microscopy images, a crucial step for downstream analysis of cellular characteristics such as morphology, cell type classification, and volumetric properties [4, 32]. However, noises in raw microscopy images can propagate into the final 3D segmentation, resulting in oversegmentation. Shown in Figure 1, this challenge occurs when individual cells are mistakenly split into multiple segments, which largely downgrades the quality of 3D segmentation results produced the current state-of-the-art methods.
Current 3D cell segmentation approaches can be broadly categorized into 2D-based and non-2D-based methods. Oversegmentation errors could appear differently depending on which category the segmentation approach falls under. The 2D-based methods reconstruct the 3D cell structure by “stitching” pre-segmented 2D slices layer-by-layer. This strategy has gained popularity recently due to its strong generalizability across diverse cell shapes and datasets, effectively eliminating the need for extensive pre-training on labeled 3D datasets. However, 2D-based approaches heavily rely on the quality of initial 2D segmentation results; thus, errors in individual slices can propagate and accumulate, resulting in oversegmentations predominantly along the imaging axis. In contrast, non-2D-based methods directly reconstruct 3D cell structures from raw microscopy images without using 2D pre-segmentated results. While this strategy mitigates some errors inherent in 2D segmentation, it requires substantial pre-training on large, dataset-specific annotations, thereby limiting its generalizability. Consequently, oversegmentation errors in non-2D-based methods typically manifest as arbitrary, irregular partitions cutting through the reconstructed cell bodies at random orientations.
To the best of our knowledge, there is currently no existing literature that explicitly addresses the nature of the oversegmentation problem. This challenge is particularly difficult because such oversegmentation errors closely resemble natural gaps separating adjacent cells.
In this work, we address the oversegmentation challenge by reformulating it as a binary classification problem, where the learning task is to distinguish oversegmentation from natural gaps between cells. We propose a geometric framework that quantitatively captures layer-to-layer 2D geometric consistency as well as overall 3D topological consistency across segmented cell bodies to reliably detect oversegmentation. In addition, we build a pre-trained classifier with the extracted geometric and topological information from flawed 3D segmentation results and then correct these identified oversegmented cell bodies through a proposed stitching method. We conduct extensive experiments across multiple plant and animal datasets to illustrate the effectiveness of our method. Our main contributions are summarized below:
(i) We formally study the oversegmentation issue in 3D cell segmentation as a specific learning problem and introduce a geometric framework for correcting oversegmentation errors. Our framework improves overall 3D segmentation quality and is compatible with various existing segmentation methods.
(ii) We introduce a novel geometry-aware metric, Geo-Wasserstein divergence, that leverages earthmover’s distance (EMD) to quantify the divergence between 2D geometric shapes of cell masks.
2 Preliminaries and Related Works
2.1 3D Cell Segmentation
The development of 3D cell segmentation methods has been greatly facilitated by deep learning architectures such as U-Net [20], which have given rise to several state-of-the-art 2D segmentation methods, including CellPose2D [17, 25, 26], Mesmer [14], and StarDist [23]. Building on these advances, 3D segmentation methods can be broadly categorized into two types: (i) 2D-based methods, which typically rely on unsupervised or weakly supervised 2D segmentation followed by subsequent 3D reconstruction, thereby benefiting from reduced complexity and strong generalizability; and (ii) Non-2D-based methods, which directly reconstruct 3D structures from raw image stacks, generally requiring fully supervised training on large labeled datasets. Within it, 2.5D-based approaches have emerged, which integrates the 2D pre-segmented information into the final 3D contextual analysis. Such hybrid framework avoids the need for large-scale 3D pretraining while leveraging the strong generalizability of 2D-based methods
CellPose [26] (2.5D-based), CellStitch [15] (2D-based), and PlantSeg [33] (non-2D-based) represent the state-of-the-art approaches in their respective categories. However, 3D-based or deep learning-based methods [8, 28, 31] such as PlantSeg often suffers from limited generalizability due to its reliance on supervised pretraining with large labeled datasets. We review the method above and provide the review for alternative approaches in Appendix A:
CellPose. CellPose is 2.5D-based approach that integrates the strengths of both 2D and 3D approaches by leveraging contextual layer awareness to develop a transferable model. It trains models to predict flow vectors for each pixel, and to generate 3D flow vectors, it averages the 2D flow vectors along the , , and planes. However, the substantial difference in sampling ratios between the plane and the -axis introduces noise, leading to oversegmentation in highly anisotropic images.
CellStitch. CellStitch is an unsupervised 2D-based approach that reconstructs 3D segmentation from CellPose2D’s pre-segmentated 2D results. It uses optimal transport to match cells between adjacent layers and introduces an interpolation method to “stitch” the discrete, layer-by-layer 2D cell masks into a continuous 3D cell body. However, the performance of such 2D-based methods heavily depends on the quality of the initial 2D segmentation. Specifically, CellPose2D would mis-segment with empty or undersegmented mask in 2D layers and causing oversegmentations in CellStitch’s 3D results.
PlantSeg. PlantSeg is a supervised non-2D-based method that employs deep learning to predict cell boundaries from 2D image stacks, followed by 3D reconstruction of cell bodies. Its segmentation quality is highly data-driven, depending on the image quality and the similarity between the input images and the training data. Specifically, PlantSeg is trained on labeled Arabidopsis thaliana ovule cells, and we empirically found that oversegmentation typically occurs in elongated cells.
2.2 Optimal Transport
Optimal transport (OT) has been applied to mutiple biological areas, including single-cell studies [22], drug perturbations [5, 9, 10], and cell alignments [11]. In our framework, OT is utilized to quantify the geometric change between 2D cell masks from adjacent layers. We introduce OT formulation along with earthmover’s distance (EMD) below.
Given measures and defined on , along with a transportation cost matrix , where quantifies the transportation cost between a pair of element , the objective is to find an optimal transport plan that determines the optimal amount of transportation between pairs which minimizes the overall transportation cost:
Eq (2.2), known as the Kantorovich formulation of optimal transport, generalizes the Monge formulation [16] by relaxing the requirement of a direct one-to-one mapping between source and target elements. For two discrete distributions and , the earthmover’s distance is defined as the minimum cost required to transform into :
| s.t. |
The EMD provides a measurement of the distance between two distributions based on the minimal work required to transform one distribution into the other. Utilizing the EMD, we introduce a novel approach to quantify the divergence between two 2D geometries: Geo-Wasserstein divergence. By approximating 2D geometries into discrete distributions of mass (see §3.2), EMD allows us to calculate the minimal transportation cost required to transform one geometry into the other. When cell masks get larger, computing the EMD becomes computationally expensive. Still, sliced Wasserstein distance [3] can be used a computationally efficient approach to estimate EMD. Typically, increasing the number of projections in sliced Wasserstein method would bring better approximation to the true EMD.
3 Main Results
3.1 Problem Formulation
We first address oversegmentation along the standard axis in 2D- and 2.5D-based models, and then generalize our approach to handle tilted cut planes at arbitrary angles. As illustrated in Figure 3, when a missing cell slice exists in layer (leaving an empty cell mask at layer ), 2D-based models would incorrectly split the entire cell into two different cells in upper section and lower section. Our task is to: (i) identify a list of suspected oversegmented candidates, consisting of upper and lower parts of cells separated by a gap; (ii) determine whether each candidate represents true oversegmentation or a natural gap, as commonly observed in loosely packed tissues like leaves, where intercellular spaces are frequent; and (iii) recover the correct 2D segmentation by predicting the cell mask for the gap layer and reconstructing the accurate 3D segmentation result. We provide a detailed explanation of these steps in the following section.
3.2 Method Pipeline
Candidates Screening. We begin by identifying suspected oversegmented candidates within the flawed 3D segmentation result. As shown in Algorithm 1, for each cell, we store its mask at the “top” (highest layer) and “bottom” (lowest layer). We then iterate over all cell pairs in the dataset. If the top-layer mask of one cell overlaps with the bottom-layer mask of another—i.e., they occupy the same position across adjacent layers—and no other cell exists between them, the pair is marked as a suspected oversegmentation candidate.
Binary Classification. For each suspected candidate, the goal is to distinguish true oversegmentations from natural gaps. We frame this as a binary classification task and train a classifier using features extracted from labeled examples of both oversegmented and natural-gap cell pairs.
Geometric Features Extraction. For each candidate pair, we quantify the geometric change at the gap layer and compare it to the typical layer-to-layer transitions within each cell. In true oversegmentations, the gap-layer change tends to be smooth and consistent with prior transitions. In contrast, an abrupt or irregular change at the gap suggests the presence of two distinct cells, indicating a natural separation. For cell masks and , we represent them as uniform distributions over their geometric regions:
where is the indicator function denoting whether lies within the boundary of , and denotes the area of mask . The Geo-Wasserstein divergence (GW divergence) between these two masks—i.e., the change in geometric shape—can then be quantified using the earthmover’s distance: . To assess whether the GW divergence between two gap-layer masks is consistent with the prior trend of mask transitions in each cell, we use Algorithm 2 to extract EMD values; each cell is indexed from layer (top) to layer (bottom). Moreover, since different cells have varying heights, the lengths of the arrays differ. To enable training, it is necessary to standardize the input length across all samples. To achieve this, we extract key statistical features from each EMD array to serve as inputs to the model. Specifically, in our pretrained model, we use summary statistics such as the median, max, min, first quartile (), and third quartile () for each array. These features, together with the gap-layer EMD value , are used as inputs to the training process.
Topological Shape Extraction. We propose a simple yet effective method from 3D topological perspective to distinguish oversegmented cell pairs from those separated by natural gaps. As illustrated in Figure 4, a key pattern observed across multiple datasets is that oversegmented cases typically exhibit a smooth shape transition at the gap layer, often following either a linear or circular progression. These patterns have biological relevance: circular shapes—wider in the middle and narrower at the ends—are characteristic of mitotic cells undergoing “circular rounding” [21, 6], while gradually tapering linear shapes are typical of structural cells such as palisade cells [18].
We quantify such linearity and circularity by analyzing the overlapping area between masks at adjacent layers. For cells exhibiting a linear pattern, the overlap area should change in a monotonically linear trend. In contrast, for circular patterns, the overlap area is expected to follow a strictly quadratic trend—characterized by a monotonous increase followed by a decrease. As shown in the Table 1, we evaluate this strict pattern across cells in the labeled datasets.
| Dataset | # of Cells | % SM |
| Filament | 5,879 | 0.963 |
| Leaf | 2,207 | 0.927 |
| Anther | 7,711 | 0.920 |
| Sepal | 6,744 | 0.908 |
| Pedicel | 7,319 | 0.869 |
| Valve | 15,817 | 0.851 |
However, by examining the cells that deviate from the strictly monotonous pattern, we observed occasional oscillations in the overlapping area changes—despite the overall shape following a clear linear or quadratic trend (see Figure 9). To improve the inclusiveness towards more candidates, we relaxed the strict constraint by fitting linear and quadratic regressions to the overlap curve and using the regression coefficient as a measure of shape consistency. For strictly monotonous cells, we assign . Each suspected candidate pair is treated as a single cell, and we apply Algorithm 3 to extract its shape information.
Classifier Training. Therefore, by leveraging both sources of information, we aim to evaluate whether the gap layer maintains geometric consistency with the preceding layers, and whether the entire cell exhibits topological consistency typical of normal cells. Our full pipeline for classifier training is illustrated in Figure 2.
We begin by constructing a training set composed of labeled true (oversegmented) and false (natural gap) cases from annotated datasets. True cases are generated by simulating oversegmentation—removing a specific layer from a 2D segmentation to create an artificial gap. False cases are obtained by applying Algorithm 1 to identify naturally separated cells. We then apply Algorithm 2 and Algorithm 3 to each candidate in the training set to extract geometric and topological shape features. To preserve all extracted information, we concatenate the features and use them as input to a multi-layer perceptron, which is trained to perform binary classification.
Cross-layer Interpolation. For final oversegmented cell pairs, we recover the mis-segmented 2D mask between them (at layer ) by interpolating between the cell mask boundaries at layers and , using the interpolation method from CellStitch [15]. The corrected 2D segmentation masks are then used to reconstruct the updated 3D segmentation. The specific implementation details of this cross-layer interpolation are provided in Appendix B.1.
3.3 Method Generalization
Recall that our original framework operates on oversegmentation results that the cutting surfaces between oversegmented cell pairs are always aligned with the horizontal plane of the standard coordinate axes. However, in non-2D-based approaches, these cutting surfaces are often tilted, irregular, and rugged. To address this, we extend our framework to effectively handle these more complex oversegmentation scenarios. We validate this generalized method on the Pancreas-B dataset, where the non-2D-based approach demonstrates better performance (see Figure 16), likely because the cells in Pancreas-B resemble those in the ovule dataset used to train PlantSeg. Notably, we observe that tilted oversegmentation is more prevalent in this dataset than the others. We also propose several recommendations to improve the end-user experience in such settings.
Candidates Screening. Since oversegmentation can occur in any direction for 3D-based approaches, we store all cell information in a graph structure and retrieve neighboring cells—i.e., those in direct contact—as suspected candidates. The number of suspected candidates in this setting is expected to be significantly larger than in the axis-aligned case described in §3.2.
Oversegmentation Cutting Surface. We define the cutting surface between two cells as the shared curved interface where the cells are in direct contact. In the previous setting, this surface was aligned with the plane, and 2D segmentation layers were sliced from the 3D volume along this standard plane. In this generalized case, we extend this concept by slicing the 3D segmentation volume layer-by-layer along a rotated reference plane—one that is aligned with the cutting surface between two oversegmented cells—thereby forming a rotated 2D segmentation representation.
Rotated 3D Segmentation Reconstruction. The first step is to estimate a reference plane from the rugged cutting surface between two oversegmented cells. We treat this cutting surface as a curvature and apply PCA to extract a best-fit plane, which serves as the estimated reference plane. Using this reference plane, we slice the 3D segmentation volume into parallel 2D layers by shifting the plane upward and downward. However, as illustrated in Step 4 of Figure 5, since all cells are stored in a discrete index format, the tilted plane intersects each layer as a series of discrete strips (aligned with integer indices), rather than forming smooth, continuous 2D geometries. To address this, we interpolate between these strips to reconstruct complete 2D cell masks. The reconstructed 2D masks are then used as input to our pretrained classifier. Specifically, we remove the masks of both cells at the reference layer (i.e., the unshifted PCA plane) and treat the remaining upward and downward segments as a pair of candidate cell fragments. The classifier then determines whether these two parts should be merged. Notably, in this setting, layer-to-layer interpolation is no longer required to correct 2D mis-segmentation—if merging is validated, we directly combine the original 3D cell bodies.
Practical Scheme. Although 2D-based and 2.5D-based approaches remain the mainstream in practice, our generalized framework aims to support a broader range of use cases for non-2D-based approaches. To provide end-users with greater control and flexibility, we introduce tunable hyperparameters into this generalized framework. Technical details of the implementation are provided in Appendix B.2.
4 Experiment
Dataset Overview. We evaluate our method on two types of datasets: publicly available, labeled plant cell datasets, and private, unlabeled animal cell datasets. Specifically, the animal datasets lack ground-truth annotation. In total, we use six different plant cell datasets, each consisting of 100 image stacks, and two animal cell datasets with 12 image stacks. Notably, each animal image stack is a large volumetric dataset containing approximately 300 image layers, with the maximum plane size reaching up to pixels. After the initial 3D segmentation, each image stack contains approximately 2,000 to 4,000 cells. Further dataset details are provided in Appendix D.
Experiment Overview. Our experimental evaluation consists of three parts: (i) synthesized test cases derived from ground-truth plant datasets, which provide deterministic labels and enable accurate evaluation using metrics such as recall and F1-score; (ii) direct augmentation of real, flawed 3D plant segmentation results, which we evaluate it across multiple state-of-the-art methods using both deterministic quantitative metrics and qualitative results validated by human experts; (iii) direct augmentation of flawed 3D segmentation results on animal datasets, which serves to assess the transferability of our method in the absence of ground truth; and (iv) ablation analysis on the Geo-Wasserstein divergence—specifically, evaluating the impact of removing layer-to-layer EMD information on the final results. We believe this experimental design and evaluation setup provide strong justification for the effectiveness and robustness of our approach.
4.1 Plant Cells: Synthesized Cases
In this task, we reserve a small portion (10 image stacks) from each plant dataset, containing true (oversegmented) and false (natural gap) test cases that are excluded from the training set. This setup allows us to evaluate the pretrained model on unseen but similar data. Since these test cases are constructed from labeled ground-truth segmentations, each has a deterministic ground-truth label, enabling precise quantitative evaluation. Results are shown in Table 2:
| Dataset | Sepal | Pedicel | Valve | Leaf | Anther | Filament |
| Recall | 0.955 | 0.854 | 0.965 | 0.988 | 0.840 | 0.854 |
| Precision | 0.963 | 0.998 | 0.983 | 1.000 | 0.985 | 0.981 |
| F1-Score | 0.959 | 0.920 | 0.974 | 0.994 | 0.907 | 0.913 |
However, the accuracy reported above serves only as a reference for the pretrained model’s performance under ideal conditions with perfect initial pre-segmented 2D results, but in pratice, pre-segmentation inaccuracies are inevitable. Therefore, we next evaluate the actual performance of our method on real, flawed 3D segmentation results generated by CellPose, CellStitch, and PlantSeg.
4.2 Plant Cells: Real Cases
In this task, we apply CellPose, CellStitch, and PlantSeg directly to the raw microscopy images to obtain the initial 3D segmentation results, which inherently contain errors due to imperfect segmented cell masks—such as shrinkage and fragmentation—commonly encountered in real-world applications. We then apply our proposed method to these flawed segmentations. This experiment is designed to evaluate the robustness of our framework in handling realistic segmentation settings.
Evaluation Metrics. We first report the quantitative results to directly assess the improvement in overall segmentation quality using standard evaluation metrics from multiple perspectives: mean Average Precision (mAP) and Jaccard Index (Jac). Specifically, the average precision (AP) is computed as , where , , and represent the numbers of true positive, false negative, and false positive masks, respectively, under a specified Intersection over Union (IoU) threshold . The mean Average Precision (mAP) is then obtained by averaging the AP scores at IoU thresholds of , , and . The Jaccard Index (Jac), by contrast, directly measures the IoU overlap between predicted and ground truth cell masks. Results are shown in Table 3.
Necessary Case-by-case Validation. However, as a new formulated problem, we would like to discuss the limitations and biases of these standard metrics in our task, and why case-by-case human evaluation remains necessary: (i) mAP may decrease even when the correction is correct. In Figure 11 (a), we show a case where two oversegmented masks were correctly stitched, yet the mAP score decreased after correction. This is because mAP and other overlap-precision-based metrics rely on a fixed IoU threshold between predicted and ground truth cell masks. During post-processing, segmentation methods often shrink the original masks, so our corrected mask at layer , built from these already shrunken masks, may fall below the IoU threshold when compared to the ground truth; (ii) Jaccard may increase even when the correction is incorrect. Although Jaccard index shows significant improvement across multiple datasets, we highlight that incorrect corrections can also inflate this metric. Shown in Figure 11 (b), the correction is actually wrong when closely compared to the ground truth masks. This issue affects all other overlapping-area-based metrics as well.
| Leaf | Sepal | Valve | Pedicel | Anther | Filament | |
| CellPose-based Results | ||||||
| mAP — CellPose | 0.4760.171 | 0.4280.133 | 0.2700.091 | 0.3580.143 | 0.3830.154 | 0.4730.172 |
| mAP — CellPoseOurs | 0.5140.179 | 0.4470.136 | 0.2860.091 | 0.3880.146 | 0.4030.151 | 0.5020.178 |
| Jac — CellPose | 0.6010.196 | 0.6110.155 | 0.5490.089 | 0.4970.174 | 0.6370.146 | 0.6460.190 |
| Jac — CellPoseOurs | 0.6550.179 | 0.6400.149 | 0.5680.087 | 0.5310.166 | 0.6640.143 | 0.7040.178 |
| (Correct, Unsure, Incorrect) | (35, 4, 5) | (142, 59, 22) | (249, 81, 43) | (136, 36, 23) | (139, 53, 34) | (100, 18, 25) |
| Case-by-case Accuracy | 0.875 | 0.866 | 0.853 | 0.855 | 0.803 | 0.800 |
| CellStitch-based Results | ||||||
| mAP — CellStitch | 0.5810.140 | 0.2680.124 | 0.4570.099 | 0.2410.181 | 0.4220.122 | 0.4020.161 |
| mAP — CellStitchOurs | 0.6060.148 | 0.2750.125 | 0.4710.100 | 0.2470.182 | 0.4380.126 | 0.4100.156 |
| Jac — CellStitch | 0.6340.136 | 0.4430.154 | 0.7200.075 | 0.3610.154 | 0.6110.163 | 0.4950.137 |
| Jac — CellStitchOurs | 0.6680.133 | 0.4560.153 | 0.7370.075 | 0.3710.155 | 0.6200.165 | 0.5040.133 |
| (Correct, Unsure, Incorrect) | (42, 38, 2) | (70, 113, 13) | (75, 82, 4) | (71, 134, 15) | (93, 141, 13) | (92, 124, 28) |
| Case-by-case Accuracy | 0.933 | 0.843 | 0.949 | 0.826 | 0.881 | 0.767 |
| PlantSeg-based Results | ||||||
| mAP — PlantSeg | 0.5640.081 | 0.3770.131 | 0.5470.098 | 0.6070.160 | 0.5630.100 | 0.4840.098 |
| mAP — PlantSegOurs | 0.5810.082 | 0.3910.132 | 0.5600.982 | 0.6210.155 | 0.5790.108 | 0.5050.097 |
| Jac — PlantSeg | 0.6920.110 | 0.5410.116 | 0.6350.079 | 0.8870.053 | 0.7480.168 | 0.6680.102 |
| Jac — PlantSegOurs | 0.7090.109 | 0.5670.123 | 0.6630.080 | 0.8900.052 | 0.7500.167 | 0.6780.112 |
| (Correct, Unsure, Incorrect) | (19, 1, 1) | (40, 2, 3) | (125, 26, 12) | (84, 12, 11) | (77, 8, 16) | (90, 11, 6) |
| Case-by-case Accuracy | 0.950 | 0.930 | 0.912 | 0.884 | 0.828 | 0.901 |
Therefore, while these quantitative metrics are important and do indicate general improvement, we emphasize that human case-by-case evaluation is still essential for a complete understanding of the correction quality. During human verification, we classify oversegmentation corrections into three categories: (i) Correct cases, where the oversegmented cell masks are present in the ground truth and belong to the same cell; (ii) Incorrect cases, where the oversegmented cell masks are present in the ground truth but actually belong to different cells (natural gap); and (iii) Unsure cases, which include ambiguous scenarios such as “hallucination”, where the correction appears geometrically plausible but lacks definitive support from the ground truth. A common source of such cases is the hallucination effect in CellPose2D (see Figure 12 and Figure 13), where highlight spots or noise in the raw image are mistakenly interpreted as valid cell masks. Although these regions are empty in the ground truth, they are treated as real masks by 3D models. Consequently, such artifacts propagate into the 3D segmentation and affect our correction process, leading to visually “correct” results that are based on flawed inputs. An example of the correct case is shown in Figure 6, where in the CellStitch row, layers and contain cell masks (highlighted using the green box) that are also appeared in the ground truth row at the same position. Our method detects these two masks and stitches them as the same cell. We provide more visualization examples of unsure and incorrect cases in Appendix C.3 and Appendix C.4.
Results Discussion. By analyzing the incorrect cases, we identified a major causes of these errors: CellPose2D often “shrinks the cell mask smaller” during 2D segmentation, especially when the cell boundaries in the raw images are blurred by a light band. This shrinkage results in the loss of geometric information, as the EMD between the two masks is also affected when one mask shrinks.
4.3 Animal Cells: Real Cases
In this task, we evaluate the transfer learning effectiveness of our pre-trained plant cell model on an unlabeled animal cell dataset (Pancreas-A, abbreviated as PA). We recommend users follow the pipeline in §3.2 to train models on their own data, but still, we offers a reference for performance on unseen and distinct cell types. Due to the lack of ground truth, we present the augmented results on CellStitch and report case-by-case verification result. Results are shown in Table 4:
| Dataset | PA-11 | PA-16 | PA-03 | PA-07 | PA-04 | PA-09 | PA-10 | PA-05 | PA-02 | PA-14 | PA-08 | Overall |
| # Correct | 9 | 10 | 19 | 29 | 21 | 25 | 35 | 52 | 28 | 52 | 94 | 374 |
| # Unsure | 1 | 3 | 3 | 5 | 3 | 9 | 8 | 7 | 8 | 21 | 24 | 92 |
| # Incorrect | 1 | 2 | 5 | 4 | 3 | 6 | 4 | 10 | 1 | 2 | 14 | 52 |
| Accuracy | 0.900 | 0.833 | 0.792 | 0.879 | 0.875 | 0.807 | 0.897 | 0.839 | 0.966 | 0.963 | 0.870 | 0.878 |
Results Discussion. Together with the results in §4.2, we observe that the performance of our framework is influenced by the resolution of the image stack and the average size of the cell masks. Higher resolutions amplify geometric details, enabling Geo-Wasserstein divergence to capture more nuanced information. We also include results on the Pancreas-B dataset in Appendix B.2, which exhibits tilted oversegmentation errors when using non-2D-based approaches such as PlantSeg.
4.4 Abation Analysis on Geo-Wasserstein Divergence
| Dataset | Leaf | Sepal | Pedicel | Anther | Filament | Valve | ||||||
| w/o EMD | w EMD | w/o EMD | w EMD | w/o EMD | w EMD | w/o EMD | w EMD | w/o EMD | w EMD | w/o EMD | w EMD | |
| # Correct | 52 | 42 | 126 | 70 | 78 | 71 | 113 | 93 | 139 | 92 | 91 | 75 |
| # Unsure | 64 | 38 | 312 | 113 | 477 | 134 | 371 | 141 | 338 | 124 | 148 | 82 |
| # Incorrect | 36 | 3 | 131 | 13 | 128 | 15 | 102 | 13 | 99 | 28 | 40 | 4 |
| Accuracy | 0.5909 | 0.9333 | 0.4903 | 0.8434 | 0.3786 | 0.8256 | 0.5256 | 0.8807 | 0.5840 | 0.7667 | 0.6945 | 0.9494 |
We trained the classifier with and without EMD features and evaluated both on CellStitch’s flawed 3D segmentation results for labeled plant cells. Case-by-case verification results are reported in Table 5. Incorporating EMD significantly improved correction accuracy by capturing geometric differences between 2D masks that area-based metrics often overlook. While 3D topological features, such as layer-to-layer overlaps, offer limited information, geometry-aware EMD reliably detects subtle layer-to-layer 2D mask shape variations critical for distinguishing true oversegmentations from natural gaps. For example, two geometries may have identical areas but entirely different shapes, then EMD could capture this shape divergence, while the area-based metrics failed. By combining 2D geometric features with 3D topological information, our framework more effectively captures geometric-topological consistency to identify oversegmentation.
5 Conclusion
In this work, we present a geometric framework to address 3D cell oversegmentation, an important yet unsolved challenge that severely affects the quality of 3D segmentation results. Our method is designed to identify and correct oversegmented regions by leveraging both 2D geometric and 3D topological information, enabling more accurate reconstruction of cellular structures. We conduct extensive experiments on both synthesized and real datasets to validate the effectiveness of our approach across in-domain and out-of-domain settings. We believe this work has meaningful implications in two areas. First, our proposed framework introduces a new direction for future research in 3D segmentation, particularly using geometry-aware methods to handle complex cellular morphologies, such as neurons and cancer cells with chaotic spatial patterns. Second, we introduce a novel metric, the Geo-Wasserstein divergence, that may inspire future work in geometry processing and low-dimensional geometric representation learning. We provide further discussion of related work in Appendix A, technical details and computational efficiency of the pipeline in Appendix B, and additional visualization results with case-specific analyses in Appendix C.
Acknowledgement
We acknowledge computing resources from Columbia University’s Shared Research Computing Facility project, which is supported by NIH Research Facility Improvement Grant 1G20RR030893-01, and associated funds from the New York State Empire State Development, Division of Science Technology and Innovation (NYSTAR) Contract C090171, both awarded April 15, 2010.
PC and YL would also like to express our appreciation to Andrew Blumberg, Abigail Hickok, Yinuo Jin for their valuable feedback and discussions.
Appendix A Further Related Works
In this section, we provide further discussion of the 2D segmentation techniques used in CellPose and CellStitch, as well as some alternative 3D segmentation approaches.
2D Segmentation Methods. We introduce the following key 2D segmentation methods that facilitate the development of 3D cell segmentation methods:
U-Net [20] is a fully convolutional neural network specifically developed for biomedical image segmentation. It employs an encoder-decoder structure with skip connections that preserve spatial information, making it ideal for cell and tissue segmentation. Hover-Net [13] combines semantic segmentation with instance-level clustering via predicted horizontal and vertical distance maps (hover maps). Hover-Net is particularly effective for segmenting densely packed cell nuclei in pathology slides. Micro-Net [19] is a lightweight CNN architecture specifically tailored for segmenting cells in resource-limited environments. It leverages depth-wise separable convolutions and inverted residuals to achieve efficient segmentation with fewer parameters. DCAN [7] simultaneously predicts pixel-wise object masks and object contours using a dual-branch CNN architecture. Effective for precisely delineating individual cells by explicitly modeling boundaries. Omni-Seg [12] uses an ensemble of CNN models trained on multiple image types to generalize across a wide range of biological imaging modalities without retraining. It applies domain adaptation and multi-task learning techniques. NuSeT [34] is a CNN model tailored specifically for nuclear segmentation in fluorescence microscopy images, combining U-Net structure with a tailored loss function to enhance segmentation accuracy in noisy conditions.
3D Segmentation Methods. We provide an overview of existing 3D segmentation techniques, along with their limitations in the context of the current state-of-the-art methods.
3DCellSeg [29] is a two-stage deep-learning pipeline designed for dense 3D cell segmentation in membrane-stained images. The pipeline is robust with only one main hyperparameter, avoiding extensive manual tuning and generalizing across different datasets and cell shapes. But still, as a supervised 3D-based method, 3DCellSeg still requires annotated training data for new cell types or imaging conditions. Its performance can degrade if applied to modalities it wasn’t trained on, necessitating retraining. Additionally, like many 3D-based methods, processing very large volumetric images can be computationally intensive comparing to 2D-based methods.
CellSeg3D [1] introduces a self-supervised 3D cell segmentation approach that eliminates the need for manual annotations. It employs a WNet3D architecture—comprising two 3D U-Nets connected sequentially—to perform segmentation by optimizing a soft normalized cuts loss directly on raw image volumes. However, as a self-supervised method, it relies on inherent image features, such as brightness differences, to delineate cells. This reliance may limit its ability to generalize effectively to end-user datasets, which often exhibit lower quality. While the performance is well on colorful fluorescent images, it may not be suitable for application in our generalized setting.
StarDist3D [30] is an instance segmentation method that represents cells (typically for nuclei) as star-convex shapes. A CNN predicts, for each pixel/voxel, the distances to the object boundary along a fixed set of radial rays, as well as an object probability score. StarDist assumes objects are approximately star-convex, so it may struggle with cells that have extremely irregular or concave shapes not well represented by a single star-shaped model. Another practical limitation is the need for training data for each new application. Due to the lack of large 3D training datasets, currently has no specific pretrained 3D StarDist models are publicly available. Still, StarDist3D achieves good performance on its specific cell types.
Go-Nuclear [27] is a recently introduced toolkit for 3D nucleus segmentation in tissue and organ datasets. Rather than a single algorithm, it encompasses a pipeline for generating training data and iteratively training models using multiple algorithms. The large curated dataset is used to train and fine-tune popular segmentation frameworks (CellPose, PlantSeg, StarDist) for nuclei in diverse organs. Yet, Go-Nuclear models are specialized for nuclear segmentation and rely on having a clearly labeled nucleus channel.
Due to the specificity of certain cell types and the lack of generalizability, as well as the absence of documentation on public open benchmarks, our work chooses to focus on CellPose, CellStitch, and PlantSeg, which have documented performance and are applicable to a wide range of cell types.
Appendix B Technical Details
B.1 Cross-layer Interpolation to Recover 2D Mis-segmented Cell Masks
As our pre-trained classifier finds the oversegmented candidates, we then need to recover the missing mask between the two oversegmented cells. To achieve this, we adopt the interpolation method proposed in CellStitch [15], which builds upon the Wasserstein interpolation framework introduced by Solomon et al. [24].
We begin by treating each matched pair of cells in consecutive slices as source and target boundaries (i.e., the contours of each cell mask). Each boundary pixel is assigned a uniform mass, and an optimal transport (OT) plan is computed to map source pixels to target pixels. For any intermediate layer, we interpolate each matched pixel pair based on the transport weights, resulting in a geometry-aware interpolation of the boundary. The interpolated boundaries are then filled to reconstruct the full cell mask at that layer. An example is illustrated in Figure 7, where a mask at the intermediate layer is generated using the cell contours from layers and .
B.2 Practical Framework to Deal with Tilted Oversegmentations
When tested with our end-users, we observed that handling tilted oversegmentation cases involved two stages of method transfer: (i) Cell type transfer, where the model transitions from in-domain plant cells to out-of-domain animal cells; and (ii) Geometric degradation, where some mask shape information is inevitably lost during the rotation and slicing process. To accommodate potential information loss, we propose a more flexible scheme that offers end-users a trade-off between precision and inclusiveness, allowing them to adjust the correction behavior based on their preference.
Penalties for Information Losses. As discussed, the reconstruction process may slightly distort the relative geometric positions of the cell masks, and interpolating between strips within each layer can alter the cross-layer overlapping areas and EMD values, thereby impacting both 3D shape indices and EMD-based features. To mitigate these information losses, we introduce a simple yet effective penalty mechanism applied to both geometric and topological features. Higher penalties suppress false positives by enforcing stricter consistency, while lower penalties allow greater flexibility, enabling the inclusion of borderline oversegmentation cases. These penalties can be tuned to match the specific objectives or tolerance levels of the end-user.
For the normalized 3D shape index retrieved from Algorithm 3, we apply a positive constant and use the penalized value as the input. For EMD information, recall that statistical features are extracted from each EMD array to ensure consistent input lengths. In our pre-trained model, we use the minimum and maximum as boundary features; here, the penalization can be adjusted by replacing the boundary features with quartiles ( and ). This approach implicitly penalizes the EMD information by narrowing the boundary features, making them stricter.
Experiment Results. We apply our framework to the 3D segmentation results generated by PlantSeg on the Pancreas-B dataset. As previously discussed, although as a 3D-based approach, PlantSeg performs relatively well on this dataset, likely due to the similar structure between cells in Pancreas-B and cells in its ovule-based training data. We conduct experiments under various penalty configurations to provide a reference for end-users. Nonetheless, we still strongly recommend that users incorporate human feedback to validate the final correction candidates, particularly if they opt to use our pre-trained model on plant cells without retraining it on their specific dataset.
| EMD Mode | Filtration | # Correct | # Unsure | # Incorrect | Accuracy | |
| 1.3 | 0.855 | 20 | 53 | 81 | 0.198 | |
| Min, Max | 2 | 0.871 | 19 | 48 | 71 | 0.211 |
| 2.5 | 0.882 | 18 | 44 | 64 | 0.220 | |
| 1.3 | 0.893 | 16 | 38 | 60 | 0.211 | |
| , | 2 | 0.920 | 14 | 25 | 47 | 0.230 |
| 2.5 | 0.941 | 11 | 17 | 35 | 0.240 |
As shown in Table 6, we adopt the human verification evaluation metric described in §4.2. Given the increased complexity as we need to handle all neighboring cells, we additionally report the filtration accuracy, which reflects the proportion of candidates excluded during screening. The penalty parameters serve as a trade-off between filtration rate (reducing human validation effort) and oversegmentation inclusion rate (capturing true positives). With appropriate penalization, only 5.9% of candidates are passed to human verification, offering end-users a flexible balance between efficiency and coverage.
B.3 Computational Efficiency
We would like to further highlight our strong applicability in terms of computation efficiency. Since the number of cells, the average size of the cells, and the density of the cells varied from datasets to datasets, we believe providing a theoretical analysis towards the computational complexity of the algorithm is impractical. Therefore, we provide direct empirical results from real applications:
Processing a large animal stack takes approximately 1 hour in total on 2 NVIDIA A40 GPUs with 48GB storage.
Running on a medium plant dataset takes around 7 minutes to identify oversegmentation candidates and around 15 minutes to stitch and recover all oversegmented cells—using CPUs only.
This speed is primarily due to:
The use of sliced Wasserstein distance for EMD computation, which is well-suited to the uniform distribution of large cell masks, speeding up the calculation through random projections.
The lightweight structure of our pre-trained MLP model.
MLP Hyperparameters. Our MLP is indeed lightweight and computationally efficient, which has 3 hidden layers with sizes 128, 64, and 32, uses ReLU activations, 0.3 dropout, a final sigmoid output.
B.4 Dealing with Multi-Oversegmentations
Although the occurrence of multiple gaps within a single cell is rare, we want to highlight that our framework is capable of handling such cases.
For example, consider a long cell body that is broken into multiple parts—say A, B, and C from top to bottom. Since our algorithm processes gaps layer by layer from top to bottom, it will first determine whether to stitch A and B. If A and B are stitched into a new cell body D, the algorithm will then assess whether to stitch D with C. This iterative process allows the framework to handle multiple fragments in a structured and consistent manner.
Appendix C Visualization Examples with Further Analysis
In this section, we present several examples of correction results produced by our framework across various datasets and viewing planes to better illustrate its effectiveness. We begin by showcasing successful correction cases, followed by a closer examination of the unsure and incorrect cases. These examples also help us highlight the limitations of deterministic metrics such as mAP and Jaccard index, which may fail to fully capture the correctness of segmentation in complex scenarios.
C.1 More Successful Examples
Figure 8 illustrates an example of correcting a mis-segmentation that is not caused by an empty mask. In the CellStitch row, the purple cell at layer and the bright yellow cell at layer remain disconnected due to an undersegmentation error produced by CellPose2D at the same position in layer . Our method successfully identifies this oversegmentation from a 3D perspective and corrects the 2D undersegmentation. The raw image clearly indicates that the three cell masks belong to the same 3D cell body. Moreover, in line 6 of Algorithm 1, we mean that there is no “complete" cell existing between the gap, with both its highest and lowest layers located between our candidates. However, we allow noisy cell masks to exist between the candidates, enabling us to also address 2D undersegmentation errors.
Figure 9 provides an example of correction viewed from the plane, illustrating how two oversegmented cells (Cell A and Cell B) are stitched along the direction. Specifically, although the raw image shows that the cell’s 3D topological shape follows a strictly circular pattern (a gradual monotonic increase followed by a decrease), 2D segmentation inaccuracies produced by CellPose2D introduce oscillations that deviate its shape from the strictly monotonic standard. However, our relaxation approach using regression alleviates this issue, as small oscillations still yield a high value, distinguishing it from natural gap cases.
Figure 10 shows an example of accurate correction in the plant cell dataset by stitching the highlighted cell masks from layer and layer . This correction is similar to the type shown in Figure 8, where we also addressed 2D undersegmentation. The true 2D segmentation in layer is recovered through cross-layer interpolation:
C.2 Inaccuracies in mAP and Jaccard Index
In Figure 11, we present two examples illustrating the limitations of mAP and Jaccard index, respectively, highlighting the necessity of human expert verification to validate the effectiveness of the corrections. The examples are drawn from CellPose-based results on plant cell datasets.
C.3 Examples for Unsure Case
Figure 12 presents an example of an unsure case caused by hallucination masks produced by CellPose2D. In the first row, the green box in the ground truth highlights an area where no cells are present. However, in the same position in the CellStitch result, two hallucinated masks (purple and yellow) are generated at layer and layer . These noisy masks lead our framework to stitch them together. While the correction appears accurate, there is no evidence from the ground truth segmentation to support our judgment. Therefore, we classify these cases as unsure, pessimistically reporting only the absolutely correct cases.
Figure 13 shows another uncertain case caused by noisy masks from CellPose2D. The green box in the ground truth marks an area without cells, but the CellStitch result includes two noisy masks (purple and orange) at layers and , which are stitched together in our result. While the correction seems accurate, the ground truth provides no evidence to confirm this, so again, we conservatively report only the fully verified cases.
C.4 Analysis on Incorrect Case
We also present example of incorrect cases to better understand the underlying issues in these situations and to provide clearer explanations for the causes of such incorrect results. As shown in Figure 14, in the row of CellStitch results, the dark blue mask at layer and the yellow mask at layer are stitched together in our results. Without the ground truth labels, this stitching appears correct, as both cells are similar in shape and located in the same position, with a missing mask in the intermediate layer. However, the ground truth row reveals that the dark blue mask is part of the large purple cell in layer , while the yellow mask belongs to the large orange cell in layer . Thus, the dark blue and yellow masks actually belong to two different cells, and their stitching is incorrect, making this an example of an incorrect case.
It is evident that both the dark blue and yellow masks only partially represent their corresponding ground truth cell masks. This highlights why we state that CellPose2D often “shrinks the cell masks smaller" than they are supposed to be. We illustrate this effect in Figure 15. When mask information is lost, both the geometric EMD measurement and the topological shape index are affected. For the EMD measurement, recall that EMD quantifies the effort required to transform one distribution into another. Losing geometric information alters the transformation, making the EMD between mask A and mask B differ from its original value. In the case shown in Figure 15, the EMD between the mis-segmented masks is smaller compared to the ground truth masks, suggesting a falsely higher similarity. For the topological shape index, which is calculated via changes in overlapping areas, as shown in Figure 14, the overlapping area between adjacent layers (e.g., layers and ) aligns more closely with the overlapping areas of preceding and succeeding layers (e.g., layers and ), as the masks shrink. These partial and altered representations lead our method to make incorrect judgments.
C.5 Comparison between PlantSeg and CellStitch on Pancreas-B Dataset
In Figure 16, we show a representative -plane slice from the final 3D segmentation results. Most of the hallucinated noisy masks generated by CellPose2D are propagated into the final CellStitch output. Although PlantSeg also exhibits oversegmentation artifacts, its overall segmentation quality is better suited for end-users’ downstream analysis.
Appendix D Dataset Information
| Dataset | # Stacks | Type | Labeled | Anisotropy () |
| Anther | 100 | Plant / Public | ✓ | |
| Filament | 100 | Plant / Public | ✓ | |
| Leaf | 100 | Plant / Public | ✓ | |
| Pedicel | 100 | Plant / Public | ✓ | |
| Sepal | 100 | Plant / Public | ✓ | |
| Valve | 100 | Plant / Public | ✓ | |
| Pancreas-A | 11 | Animal / Private | ||
| Pancreas-B | 1 | Animal / Private |
Table 7 shows the information for both plant and animal dataset. All the plant-type datasets are publicly availiable at the following site [2]. For plant-type data, the voxel resolution is unknown. For animal-type data, the voxel resolution () is .
Pre-training Setup. For CellPose, we use the pre-trained cyto2 model to generate 2D pre-segmented results. We also tune the diameter case-by-case, from 60 to 100, based on the cells in each datasets. For PlantSeg, we use the pre-trained generic_confocal_3D_unet model to generate 3D segmentation results.
Appendix E Future Works and Limitations
Labeled Dataset for Pre-training. As discussed in previous sections, multi-stage transfer learning and geometric information loss can reduce the robustness of the correction framework. Constructing a pre-trained classifier tailored to the target dataset would significantly mitigate this issue. However, doing so typically requires substantial human-labeled data. To address this, future work could incorporate active learning strategies to automatically identify the most uncertain or influential oversegmentation cases for annotation. This approach would reduce labeling effort while enhancing generalization across diverse datasets.
Tackle Undersegmentations. Future work could also study how to extend the framework to also detect and correct undersegmentation (merging of multiple cells) and fragmentation (non-contiguous segments of a single cell), creating a more comprehensive post-processing tool with our work.
Broader Applications of Geo-Wasserstein Divergence. As a novel metric for capturing 2D geometric divergence, Geo-Wasserstein has shown promise in quantifying differences between cell masks. However, its potential remains underexplored. Future work could investigate its broader applications in other geometric domains, such as shape evolution, medical imaging, or morphometrics. Additionally, assessing its robustness across a wider range of cell types would be valuable. Due to the scarcity of publicly labeled 3D datasets, our evaluation was limited to plant and animal cells commonly used by end-users.